# pyCFS-data: Data Processing Framework in Python for openCFS

### v2

@@ 2.2 Writer class (CFSWriter)
[…]
The Quantity attribute can be chosen as any string in general. However, if the exported data will be the input of a subsequent openCFS simulation, openCFS variable names must be used for the declaration of field quantities. Thus, for the acoustic PDE, one of the following names must be chosen.
General acoustic and fluid mechanic quantities:
  • acouPressure
  • acouVelocity
  • acouPotential
  • acoutIntensity
  • fluidMechVelocity
  • meanFluidMechVelocity
  • fluidMechPressure
  • fluidMechDensity
  • fluidMechVorticity
  • fluidMechGradPressure
Aeroacoustic Source Terms:
  • acouRhsLoad (general)
  • acouRhsLoadP (Perturbed convective wave equation)
  • vortexRhsLoad (Vortex sound theory)
  • acouDivLighthillTensor (Lighthill’s acoustic analogy)
[…]
@@ 3.3 Synthetic Noise Generation and Radiation (SNGR) source computation (sngr)
Computation of synthetic flow acoustic sources based on the theory of C. Bailly [1] for the SNGR similar to the implementation described in [5].
@@ 4.1 Ansys Mechanical (ansys_io)
[…]
@@ 4.7 STL reader (stl_io)
5 Acknowledge

### v3

@@ 2.2 Writer class (CFSWriter)
  The Quantity attribute can be chosen as any string in general. However, if the exported data will be the input of a subsequent openCFS simulation, openCFS variable names must be used for the declaration of field quantities. Thus, for the acoustic PDE, one of the following names must be chosen. General acoustic and fluid mechanic quantities: acouPressure acouVelocity acouPotential acoutIntensity fluidMechVelocity meanFluidMechVelocity fluidMechPressure fluidMechDensity fluidMechVorticity fluidMechGradPressure Aeroacoustic Source Terms: acouRhsLoad (general) acouRhsLoadP (Perturbed convective wave equation) vortexRhsLoad (Vortex sound theory) acouDivLighthillTensor (Lighthill’s acoustic analogy)
CFSResultInfo: Data structure containing result information for one result. Its attributes and properties are depicted in Fig. 6.
  The CFSResultInfo structure is redundant to the metadata contained within a CFSResultArray object!
@@ 3.1.4 Projection-based interpolation
3.1.5 Radial basis function interpolation and differentiation
This module provides functions to interpolate data and compute gradients using radial basis functions (RBF). Implemented kernel functions are: Gaussian RBF kernel, Multiquadratic RBF kernel, and Wendland C2 RBF kernel.
The implemented methods include interpolation and gradient computation of a field quantity using global or local RBF approximation. Parameters include
  • kernel: Type of RBF kernel to use. Available kernels include gaussian, multiquadratic, and wendland_c2.
  • epsilon: Shape parameter for the used RBF kernel function
  • smoothing: Smoothing parameter for the kernel matrix
for the global interpolation and additionally for the local interpolation
  • neighbors: Number of nearest neighbors to use for local interpolation
  • min_neighbors: Minimum number of neighbors to use based on a local neighborhood radius $r=f_{r}\frac{1}{N}\sum_{i=1}^{N}d_{i}$ with factor $f_{r}$, and the distance of the nearest neighbor distances $d_{i}$
  • radius_factor: Factor $f_{r}$ to determine the local neighborhood radius.
@@ 3.3 Synthetic Noise Generation and Radiation (SNGR) source computation (sngr)
Computation of synthetic flow acoustic sources based on the theory of [1] for the SNGR with additional description found in [13]. The implementation is described in detail in [16].
@@ 3.4.3 Fit geometry (fit_mesh)
3.5 Spatial derivatives (derivatives)
Radial basis function based evaluation of spatial derivatives (see section˜3.1.5).
3.6 Time signal processing
Module containing methods for time domain data processing operations.
3.6.1 Time derivative (timeproc.time_derivative)
Compute the time derivative of a result array using a noise-robust differentiator. The time derivative of the desired quantity $\partial q/\partial t$ is computed by a smooth noise-robust differentiator, which supresses high frequencies and is precise on low frequencies, according to [3]. For an efficient and robust calculation the order of the differentiator is set to $N=5$ and the time derivative is calculated by
$\frac{\partial q}{\partial t}\approx\frac{2(q_{1}-q_{-1})+q_{2}-q_{-2}}{8\Delta t}\,$, (5)
where the index of $q$ defines the time step relative to the time step of which the derivative is calculated and $\Delta t$ is the time step size.
Parameters include
  • boundary_treatment: The start and end of the time series can either be removed (remove), left untreated ((None)), or treated by one-sided estimators according to [4] (one-sided).
3.6.2 Field fft (field_fft)
Apply Fast Fourier Transformation (FFT) to the field data along a specified axis.
3.6.3 Dynamic mode decomposition (DMD) (field_fft)
Apply Dynamic Mode Decomposition (DMD) to the field data.
  This submodule make use of pydmd, which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[dmd].
@@ 4.1 Ansys Mechanical (ansys_io)
  This submodule make use of Ansys Data Processing Framework (DPF), which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[ansys].
  This submodule make use of Ansys Data Processing Framework (DPF), which is in Beta development stage. Therefore, compatibility might break rapidly.
  The methods in this submodule make use of Ansys Data Processing Framework (DPF), which requires a licensed installation of the according DPF Server!
@@ 4.4 Exodus file format (exodus_io)
4.5 MeshIO interface (meshio_io)
Load meshes via meshio into PyCFS’s CFSMeshData format.
  This submodule make use of meshio, which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[meshio].
@@ 4.8 STL reader (stl_io)
5 Applications
Recently, pyCFS-data was used in hybrid experimental-numerical data processing routines to estimate sound power radiation of shell and plate structures including a Gaussian Process Regression [15]. Furthermore, pyCFS-data is heavily used in the pre-processing of aeroacoustic source terms of the aeroacoustic wave equation based on Pierce operator [7, 11, 8] and the compressible perturbed convective wave equation [12, 6]. It is a Python alternative of the develope C++ alternative openCFS-data [10]
Acknowledge
@@ 5 Applications
Funding information
A.W. and S. S. received funding from the COMET project ECHODA ("Energy Efficient Cooling and Heating of Domestic Appliances") under FFG project No. FO999913905. ECHODA is funded within the framework of "COMET - Competence Centers for Excellent Technologies" by BMIMI, BMAW, the province of Styria, and SFG. The COMET program is managed by FFG. S. S. received funding from the FFG under project No. FO999913972.